OpenStudy (anonymous):

∠PQR and ∠ABC are complementary angles and ∠PQR is eight times as large as ∠ABC. Determine the measure of each angle. Students must show all work step by step

OpenStudy (anonymous):

here is what i got for it can someone take a look at it for me and let me no if it's good

OpenStudy (anonymous):

Let ABC = x Then PQR = 90-x Equation: (90-x) = 8x 90 = 9x x = ABC = 10 degrees Then 90-x = PQR = 80 degrees

OpenStudy (anonymous):

girl i be struggling with geometry questions like this myself ppl barely answer them

OpenStudy (anonymous):

lol thank you I gues il just have to go with what i have

OpenStudy (anonymous):

or u can ask the question yahoo answers

OpenStudy (anonymous):

*on

OpenStudy (anonymous):

oh i have that thanks

OpenStudy (anonymous):

k